OpenStudy (anonymous):

A hat contains 22 names, 10 of which are female. If five names are randomly drawn from the hat, what is the probability that at least four female names are drawn?

OpenStudy (anonymous):

420

OpenStudy (anonymous):

YA KONT

OpenStudy (anonymous):

The answer is actually 0.105 YA DUNCE

OpenStudy (anonymous):

Now, can someone actually help me solve this step by step

OpenStudy (anonymous):

10 females 12 males 5 will be randomly be chosen. Restriction: at least 4 females should be chosen (it can be 4 F or 5 F) This is equals to only 1 OR none males. So it will be: (10 C 4 x 12C 1) + (10 C 5 x 12C 0) Without restriction, there are 22 people and 5 will be chosen, so 22C5 Probability of having at least 4 females will be: \(\sf \Large \frac{( _{10} C_4 \times  _{12}C_1) + ( _{10} C_ 5 \times  _{12}C_0)}{ _{22}C_5}\)

OpenStudy (anonymous):

Thank you so much! I have been waiting for someone to help me with this problem for a while and havent gotten much of a response. Now thanks to you I understand this problem easily :)